OpenStudy (anonymous):

Determine the vertex coordinates for: h(x) = (2x - 5)(2x + 3)

OpenStudy (anonymous):

Expand the right side: y = 4x^2 - 4x - 15. The x-coordinate of the vertes is -(-4/(2*4)) = 1/2. f(1/2) = (1 - 5)(1 + 3) = -16. The vertex is at (1/2, -16).

OpenStudy (anonymous):

tahnks